OpenStudy (anonymous):

Ace's grandfather is 3 years older than Aces grandmother. Their combined age is 121 years. How old is each?

OpenStudy (alexandervonhumboldt2):

f=m+3 f+m=121 solve this sytem of equations: as f=m+3 we can substitute this instead of f into second equation: m+3+m=121 m=59 grandmother is 59 grandfather is f=m+3=59+3=62 62+59=121 grandfather is 62 grandmother is 59